OpenStudy (studyboiiiiiii):

Solve for x: 4 - (x + 2) < -3(x + 4)

Awolflover1 (awolflover1):

add 3x to both sides \[−x+2+3x<−3x−12+3x\]

Awolflover1 (awolflover1):

\[2x+2<−12\]

Awolflover1 (awolflover1):

Then Subtract 2 from both sides. \[2x+2−2<−12−2\]

Awolflover1 (awolflover1):

\[2x<−14\]

Awolflover1 (awolflover1):

Now Divide both sides by 2 \[\frac{ 2x }{ 2}<\frac{ -14 }{ 2}\]

OpenStudy (redheadangel):

Would one of y'all help me? Given the equation, y = 3(0.87)^x , where y is the amount of money in thousands in x years, describe what is happening? The amount of money is decreasing at a rate of 13% per year The amount of money is decreasing at a rate of 87% per year The amount of money is decreasing at a rate of 13 per year The amount of money is decreasing at a rate of 87 per year

alones (alones):

@redheadangel make a post.

OpenStudy (redheadangel):

Its not letting me:(

OpenStudy (redheadangel):

There's no like "post button" Only the box

OpenStudy (redheadangel):

@AloneS

Awolflover1 (awolflover1):

do you have grammar thing on ?

OpenStudy (redheadangel):

yeah, Grammarly? @Awolflover1

Awolflover1 (awolflover1):

yes you need to turn it off Grammarly doesn't work with os

OpenStudy (redheadangel):

oh, is that why its not letting me post? @Awolflover1

alones (alones):

@redheadangel i posted it for you (:< Enjoy.

Awolflover1 (awolflover1):

yes red

OpenStudy (redheadangel):

oh thank you! @Awolflover1

Awolflover1 (awolflover1):

you're welcome ^_^